OpenStudy (anonymous):

5 workers can produce 250 cakes in 6 hours. If one of the workers falls ill, how long would it take the rest of the workers to produce 400 cakes? How-to please

Parth (parthkohli):

5 workers take \(\large \frac{6}{250} = \frac{3}{125}\) hours to bake one cake. 1 worker takes \(\large \frac{3\times 5}{125} = \frac{15}{125} = \frac{3}{25}\) hours to bake a cake. 4 workers take \(\large \frac{3 }{25\times 4} = \frac{3}{100}\) hours to bake a cake. So, to bake 400 cakes, they will take \({\large \frac{3}{100}} {\times } 400\) hours.

OpenStudy (anonymous):

Thanks :D

Parth (parthkohli):

No problem!